A company has borrowed K800,000 from a bank. The loan is to be repaid by level instalments, payable annually in arrears for 10 years from the date the loan is made. The annual repayments are based on an effective rate of interest of 8% per annum. I. Calculate the amount of the level annual payment which will be paid over the 10-year term.

Construct an amortization schedule showing the capital and interest components over the first four years of the loan ?


1
Expert's answer
2022-03-02T10:27:24-0500

`1)

"A=800000(0.08+""\frac{0.08}{(1+0.08)^{10}-1}" ")" "=119223.59"

2)